OpenStudy (aravindg):

when do we say that 2 circles intersect ortogonally?

OpenStudy (apoorvk):

There is this general criteria that for the standard equation of two circles S_1 and S_2 \[(Standard ~equation = x^2 + y^2 + 2gx + 2fy +c = 0)\] So, for orthogonal circles, \[\large 2(g_1g_2 + f_1f_2) = c_1 + c_2\]

OpenStudy (aravindg):

can u say what we mean by orthogonal intersection ?

OpenStudy (apoorvk):

Orthogonal intersection means they intersect at right angles - rather the tangents to both the circles at their points of intersection are perpendicular. Like this:

OpenStudy (aravindg):

so when all is orthogonal intersection possible?is it there when the circles just touch each other?

OpenStudy (apoorvk):

I'm sorry, I thought you were looking for the mathematical formula, initially. :P

OpenStudy (aravindg):

no prob :) u see the formula was there in my material but theory was nt xplained

OpenStudy (apoorvk):

No - try to think! When they just 'touch' they are tangentially to one another - so there is no chance of them being orthogonal at all. Like I said - the mathematical relation above relates if two circles intersect orthogonally.

OpenStudy (apoorvk):

Hmm, getting a 'feel' of the theory is essential ;)

OpenStudy (aravindg):

what we mean by chord of contact?

OpenStudy (aravindg):

 do these 4 lines form a square or rectangle ?